Write an equation in point-slope form for the line that has a slope of 6 and contains the point (-8,-7).
Answer by TimothyLamb(4379)

point-slope form:
y - y1 = m(x - x1)
---
y + 7 = 6(x + 8)
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php